can u please help me solve this problem: two positive integers are in the ratio 1:3. if their sum is added to their product the result is 224. find the integers.
Answer by scott8148(6628)

let x=first, so 3x=second ... 3x^2+4x=224 ... 3x^2+4x-224=0 ... (3x+28)(x-8)=0